OpenStudy (anonymous):

how do i find the indefinite integral of 2x^9-6x^2+3/x^9

OpenStudy (anonymous):

is that all over x^9?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the i would suggest simplifying it first , like for the first term it would be 2x^9/x^9=2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so, your good, or you still have some doubts

OpenStudy (amistre64):

lagrange did good on this, so what is the question that remains?

OpenStudy (anonymous):

what

OpenStudy (amistre64):

recall the power rule for integrations\[\int x^ndx=\frac{1}{n+1}x^{n+1}\] thats the only thing that would be needed to finish it up

OpenStudy (anonymous):

i really dont know how to start the problem

OpenStudy (inkyvoyd):

Hint

OpenStudy (inkyvoyd):

1/(x^n)=x^(-n)

OpenStudy (amistre64):

you start it with algebra ..... simplify the integrand

OpenStudy (inkyvoyd):

THe power rul for integration still applies, unless n=-1

OpenStudy (anonymous):

so 2x^10/x^10?

OpenStudy (inkyvoyd):

Remember that you can seperate each term with the distributive property.

OpenStudy (anonymous):

how

OpenStudy (inkyvoyd):

(2x^9-6x^2+3)/(x^9) =(2x^9)/(x^9) -(6x^2)/(x^9) +3/(x^9)

OpenStudy (inkyvoyd):

Just pre algebra, my friend.